OpenStudy (anonymous):

Write a polynomial function with rational coefficients so that P(x)=0 has the given roots. #16. 2 + i and 1-√5 #15. 3i and √6

OpenStudy (anonymous):

ick

OpenStudy (anonymous):

i take it these are two separate problems right?

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

which one you want to do first? #15 is the easier of the two

OpenStudy (anonymous):

Lets tackle that one first then.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the conjugate of \(3i\) is \(-3i\) and the quadratic with those two roots is \[(x-3i)(x+3i)=x^2+9\] quick and easy

OpenStudy (anonymous):

the conjugate of \(\sqrt6\) is \(-\sqrt6\) and the quadratic with those zeros is \[(x-\sqrt6)(x+\sqrt6)=x^2-6\]

OpenStudy (anonymous):

your job is now to multiply out \[(x^2+9)(x^2-6)\] without making an algebra mistake you good for that one?

OpenStudy (anonymous):

Eh..I'm stuck on multiplying the squares

OpenStudy (anonymous):

\(x^2\times x^2=x^4\)

OpenStudy (anonymous):

you get \[x^4-6x^2+9x^2-54\] or \[x^4+3x^2-54\]

OpenStudy (anonymous):

ok for that one? the second one is harder

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

you need a quadratic with \(2+i\) as a zero

OpenStudy (anonymous):

there are several ways to find it i will show you an easy one, then a really really easy one but the really really easy one requires memorizing something

OpenStudy (anonymous):

Thats fine

OpenStudy (anonymous):

work backwards \[x=2+i\\ x-2=i\] then square both sides and get \[(x-2)^2=i^2\] or \[x^2-4x+4=-1\] add \(1\) and get \[x^2-4x+5\] as your quadratic

OpenStudy (anonymous):

the really really easy way is to know that if \(a+bi\) is a zero, then the quadratic is \[x^2-2ax+(a^2+b^2)\] in your case you have \(2+i\) so \(a=2,b=1\) and the quadratic is \[x^2-2\times 2x+(2^2+1^2)=x^2-4x+5\]

OpenStudy (anonymous):

as for \(1+\sqrt5\) without the \(i\) it is \[x^2-2x+(1^2-\sqrt5)^2=x^2-2x-4\]

OpenStudy (anonymous):

your last job is to multiply out \[(x^2-4x+5)(x^2-2x-4)\] i would cheat

OpenStudy (anonymous):

Okay here's another thing I'm having trouble w/ ... Finding all rational roots for P(x)=0 #26. P(x)=x^3-x^2-34x-56

OpenStudy (anonymous):

cheat

OpenStudy (anonymous):

does it say "find the rational roots" or "find all POSSIBLE rational roots"?

OpenStudy (anonymous):

here is the solution http://www.wolframalpha.com/input/?i=x^3-x^2-34x-56

OpenStudy (anonymous):

It just says "Find all rational roots"

OpenStudy (anonymous):

open the link, there they are

OpenStudy (anonymous):

It doesn't show me how it found the roots. Can you show me how to do one so I can do the rest?

OpenStudy (anonymous):

P(x)=3x^3+x^2-8x-12